OpenStudy (anonymous):

Johns weekly gross pay is 250, but 23% of his check is withheld as deductions. How much of his check is withheld?

OpenStudy (anonymous):

$250 * 23% = $250*.23 = $57.5

OpenStudy (anonymous):

250 * 23% = 250 * 23/100 = 250 * 0.23 = 57.5